If x-bar = 23 ounces, σ = 1.0 ounces, and n = 16, what will be the ± 3σ control limits for the x-bar chart?
1
Expert's answer
2020-12-02T07:28:38-0500

The process average is 23

 This means that σ is 0..4 ounces


σx will be computed as σ/sqrt(n) where n= represents the size of the sample


σx=0.4/sqrt(16)=0.4/4= which gives 0.1


UCL=Process average when added to zσx  LCL will translate to Process average of -zσx


UCL=23+3x0.1 which gives 23.3


LCL will be computed by 23-3X0.1 will give 22.7